### v1

@@ 3.1 Video Selection
In constructing the dataset, we adhered to three principles to address the challenges in video understanding capabilities: our dataset, FunQA, is visual centered and emphasizes counter-intuitive reasoning, spatial-temporal reasoning. Based on these principles, we collect 4365 videos from 3 different art genres and created three subsets: HumorQA, CreativeQA, and MagicQA.
@@
3.2 Task Definition
Detailed Description Task The description task aims to evaluate the model’s information extraction capabilities, serving as a fundamental aspect of video understanding. Across all three subsets, this task requires providing a free-text answer that describes the selected moment. Furthermore, this task allows for analysis of how the model extracts information and generates answers for subsequent tasks. By examining the model’s performance in this task, we gain insights into its ability to extract relevant information and generate meaningful responses.
Higher Level Tasks In addition to the three main tasks, we design higher-level tasks to enhance the model’s inference abilities on counter-intuitive videos. Title Task in HumorQA and CreativeQA requires generating a concise title summarizing the video’s content. Creative Scoring Task in CreativeQA involves rating the creativity of videos between 0 and 20. Magic Method Task in MagicQA requires the model to explain clearly the rationale behind the magic, and its purpose is to test the model’s ability to reason more deeply. To ensure the accuracy of the answers, this task is only partially annotated and appears only in the test set, details of which can be found in Appendix A.1.
In addition, some visual QA benchmarks [37] have begun to adopt free-text answers; this flexible answer format enables multimodal LLMs to go beyond multiple choice options (Large Language Models) [35, 36]; for Video QA, however, free-text are less common. To bridge this gap, we also consider some free-text answer evaluations, demanding enhanced video comprehension capabilities from the model. It will also assess the model’s ability to generate reasonably lengthy textual responses.
3.3 Dataset Statistics
FunQA contains 4,365 counter-intuitive video clips and 311,950 question-answer pairs, the total length of these videos is 23.9h and the average length of video clips is 19 seconds. FunQA consists of three fine-grained subsets, each one containing well-designed tasks. The specific numbers and split of videos can be seen in Figure 2 (d). Each subset’s video lengths can be seen in Figure 2 (d). The specific number of QA pairs for each task can be seen in Figure 2 (e).
For our localization task, the timestamp heat map for the three different types of videos can be seen in Figure 2 (g), which shows the high-frequency time span of the answer. For the description and reasoning tasks, the average length of the words in their free-text answers reached 34.24, which is much longer than existing VideoQA datasets (e.g., 8.7 in Activity-QA [50], 11.6 in NExT-QA [18]). The specific word count of each task is shown in Figure 2 (h). FunQA has a well-established annotation process and high annotation quality, the result of our annotation consensus evaluation can be seen in Figure 2 (i). For each video category, more than 90% of the annotations exhibit a high level of consensus, with only 1% of the content showing low consensus. Approximately 8% of the data demonstrates variations in consensus, thus highlighting the objectivity of our dataset.
HumorQA, CreativeQA, and MagicQA word clouds are shown in Figure 2 (a-c). More statistics and FunQA full word cloud are given in Appendix A.3.
3.4 Dataset Construction Pipeline
FunQA dataset construction pipeline was in three stages: Pre-processing, Manual Annotation, and Post-Processing. The whole process took about 900 hours with over 50 highly educated undergraduates as part-time annotators, and we paid crowd-workers a target of ¥25/hr. More details of the dataset construction pipeline can be seen in Appendix A.1.
Pre-processing Initially, we crawled videos from TikTok, Bilibili, and YouTube (these videos are a collection of surprising videos). Then we performed a two-stage manual cleaning and cutting process on the collected videos to ensure counter-intuitive features and video quality and to exclude non-ethical content and sensitive information, resulting in a cut video clip.
Manual Annotation We annotated the videos according to the characteristics of different tasks design in Chinese. We screen and train the annotators to ensure the accuracy and high quality of the annotation, and finally produce the original annotated files. After the first round of annotation, we conducted a secondary annotation of 10% of the tasks and performed Consensus Evaluation to ensure the objectivity of our annotations.
Post-processing Based on our carefully designed tasks and high-quality annotations, we expanded our dataset using GPT-3.5. Firstly, we automatically translated the Chinese annotations into English. Subsequently, we generated more QA pairs that were faithful to the original ideas but presented differently. This not only made FunQA multilingual but also increased the number of QA pairs to 312K. Additionally, we created diverse task types, such as FunQA multiple-choice and FunQA dialogue. More details are given in Appendix A.2 and Appendix B.
4 Limitations and Future Work
This paper has two limitations. 1) Current FunQA dataset mainly includes video-level data and annotations, but denser annotations (e.g., like PVSG [68]) can be developed to explore more possibilities of video reasoning. Examples include detailed spatial and temporal annotations, such as captions corresponding with specific time axes and annotations of object level. 2) The raw annotations are completed by the annotator in Chinese. In the process of translating into English, we first use GPT to polish and supplement Chinese annotations, making the text as thorough as possible, as short or incomplete text may result in misunderstanding. However, there may still be differences due to cultural diversities between the two languages.
In the future, we expect to expand our dataset with denser and more diverse annotations. Also, new metrics will be explored to better evaluate models’ performance, especially in open-ended questions which lack in-depth metrics. Finally, we hope to provide directions for advancing models toward deeper video reasoning. Specific ideas include that Otter can sample more frames from the input video to improve causal reasoning capacity. Furthermore, in the future, when refining tuning datasets such as MIMIC-IT [69] for visual language models, it may be beneficial to incorporate the FunQA training set. However, to ensure fair comparisons and prevent data leakage, it is advised that future research does not utilize the FunQA testing set.
Acknowledgments and Disclosure of Funding
Appendix A More details of FunQA dataset
A.1 Dataset Construction Pipeline
@@
Figure 3: Example of Magic Method Task. During the annotation process, we discovered that even as humans, it is difficult to fully understand the complete principles behind the implementation of magic tricks right from the beginning.
@@
Figure 4: FunQA Video Selection and Clip Principle. We have a zero-tolerance policy regarding the inclusion of offensive content in our dataset. During the video sourcing process (video selection and video clipping stages), we ensure that such content is completely eliminated.
@@
Figure 5: Examples of incorrect annotations and the suggested modifications. Here we clarify the meanings of different cell colors. Red: Errors made by the annotator during the annotation process. Yellow: Types of annotation errors. Green: Reasons for the errors and suggested corrections.
@@
Figure 6: GPT prompt for Translation and augmentation. Under our carefully designed prompt, GPT-3.5 can automatically generate English translations and five synonymous sentences from our high-quality Chinese annotations, thereby expanding FunQA dataset.
@@
Specifically for the Magic Method Task (M4), we found that most annotators were not professional magicians, and even when watching instructional magic videos, it was challenging for them to provide complete and reasonable annotations for magic methods. Therefore, for M4, we only provided a small number of examples, which do not exist in the FunQA dataset. Figure 3 shows an example of M4.
We conducted strict quality reviews in real-time for the annotations, including but not limited to checking if the annotations meet the requirements, if the selected videos contain sensitive content, and if the annotation texts contain sensitive content. The annotation principles (video selection and video clip, etc) can be seen in Figure 4. Figure 5 also illustrates some errors encountered during the annotation process and the suggested modifications provided. After completing all the annotations, we conducted a consensus evaluation study to assess the objectivity of our annotations. We randomly selected 10% of the videos and asked annotators to provide a consistency score (high consensus, medium consensus, low consensus) between their first and second annotations, considering all the previous annotations shown to them.
Copyright & License We respect the copyright of each video. For content downloaded from streaming platforms, we obtained permission from the respective uploaders. For Kasou Taisho, we submitted a fair use request to the official website of Tokyo TV in Japan.
@@
A.2 Language Augmentation
For each QA pair, we first asked GPT-3.5 to translate the previous Chinese answers into English, and then let GPT-3.5 give five separate answers with the same meaning but different linguistic expressions. Afterward, we filtered out incorrect and incomplete generations. Figure 6 shows the prompt we gave to GPT-3.5.
However, as mentioned in the main text in the limitation, the translation of GPT-3.5 will be ambiguous when the original text uses Chinese harmonics and some special Chinese meanings, and this problem mostly occurs in the Title Task, we screened such problems by hand and modified them, as shown in Figure 7.
Figure 7: Issues in translation caused by Chinese-English cultural differences. In this example of translating and expanding a humorous video title, GPT-3.5 failed to understand the meaning of the original Chinese title, and we filtered out such data and made manual changes.
A.3 Data Statistics
We collected raw videos from multiple video platforms, including short videos, long videos, and video clips, and the detailed data can be seen in Table 2. The word cloud shaped by FunQA logo of the all annotation word is shown in Figure 8.
Table 2: Statistics of the FunQA raw data.
Raw Data Type Source Videos Avg.len(s) Total.len(h)
Humor Daily Life (Human) Bilibili 351 182 15.8
TikTok 1296 14 5.18
Nature (Animal) Bilibili 230 133 8.52
Creative Performance Kasou Taishou 26 6060 43.77
@@ A.3 Data Statistics
Figure 8: Word cloud of FunQA free-text answer.
Appendix B FunQA Extension Datasets
@@
Figure 9: GPT prompt for FunQA Multi-choice. We give GPT-3.5 prompt as system, and then enter the description and explanation of the video, and GPT-3.5 will output the data in the form of multiple-choice questions according to the prompt.
@@
Figure 10: GPT prompt for FunQA Dialog. We give GPT prompt as system, and then input the description and explanation of the video, and GPT-3.5 will output the data in the form of dialogue according to the prompt.
B.1 FunQA Multi-choice Dataset
@@
FunQA Multi-choice Dataset is prepared to provide training and testing for arbitrary models, in this dataset our QA pairs are in the form of multiple choice, the answer is a word, phrase, or short sentence, and the type of questions are all descriptions. An example of the prompt given to GPT-3.5 and the data generated can be seen in Figure 9.
B.2 FunQA Dialog Dataset
@@
Most of the current LLMs are in the form of dialogues. To cater to their data input, we produced the FunQA Dialog dataset, in which we used GPT-3.5 to convert QA pairs into recursive dialogues with added context. In addition to this, we also let GPT-3.5 freely generate future dialogue development, which greatly expands the amount of information. An example of the prompt given to GPT-3.5 and the data generated can be seen in Figure 10.
Appendix C More Details of Experiment
C.1 Significance of New GPT-4 Based Metric

### v2

@@ 0.A.1 Dataset Construction Pipeline
Video Selection In constructing the dataset, we adhered to three principles to address the challenges in video understanding capabilities: our dataset, FunQA, is visual centered and emphasizes counter-intuitive reasoning, spatial-temporal reasoning. Based on these principles, we collect 4365 videos from 3 different art genres and created three subsets: HumorQA, CreativeQA, and MagicQA.
@@
Table A1: Comparison between FunQA and other existing benchmarks (Complete Version). Compared to other datasets, FunQA revolves around the captivating realm of interesting and counter-intuitive videos. The tasks within FunQA are specifically designed to challenge the vision capabilities of models, requiring strong skills in producing an in-depth description, interpretation, and spatial-temporal reasoning. Here we clarify the abbreviation in the table. For annotation type: [Uncaptioned image] denotes Manual Annotation and [Uncaptioned image] for Automatic Annotation; Avg Len denotes video average length; # Clips means number of video clips; VC for visual-centric, Des. for Description, Exp. for Explanation, STR for Spatial-temporal Reasoning, MC means Multiple Choice QA, and OE shows Open Ended QA with Average Word Count per response.
Dataset Domain [Uncaptioned image] or [Uncaptioned image] Video Question Answer
Avg Len # Clips # QA VC Des. Exp. STR MC OE
TGIF-QA [20] Social Media [Uncaptioned image] 3s 72K 165K 2.1
MSRVTT-QA [60] Social Media [Uncaptioned image] 15s 10K 244K 1.0
ActivityNet-QA [68] Social Media [Uncaptioned image] 180s 6K 58K 1.9
MSVD-QA [60] Social Media [Uncaptioned image] 10s 2K 51K 1.0
YouTube2Text-QA [66] Social Media [Uncaptioned image] 10s 10K 123K N/A
AGQA [16] Social Media [Uncaptioned image] 30K 10s 192K TBD
AVQA [64] Social Media [Uncaptioned image] 60s 9K 57K N/A
NExT-QA [59] Daily life [Uncaptioned image] 44s 5K 52K 2.6
Social-IQ [69] Daily life [Uncaptioned image] 99s 1K 8K N/A
STAR [58] Daily life [Uncaptioned image] - 23K 60K N/A
FIBER [6] Daily life [Uncaptioned image] 10s 28K 2K TBD
MovieQA [54] TV shows [Uncaptioned image] 203s 7K 6K N/A
TVQA [32] TV shows [Uncaptioned image] 76s 22K 153K N/A
TVQA+ [33] TV shows [Uncaptioned image] 8s 4K 30K N/A
KnowIT-VQA [14] TV shows [Uncaptioned image] 60s 12K 24K N/A
SUTD-TrafficQA [61] Traffic [Uncaptioned image] 5s 10K 623K N/A
MarioQA [44] Games [Uncaptioned image] 5s 188K 188K 2.0
CLEVRER [67] Synthetic Videos [Uncaptioned image] 5s 20K 305K N/A
Env-QA [12] Egocentric [Uncaptioned image] 20s 23K 85K N/A
HumorQA (Ours) Daily life [Uncaptioned image] 7s 2K 141K 28.2
CreativeQA (Ours) Performance [Uncaptioned image] 48s 1K 79K 59.1
MagicQA (Ours) Magic shows [Uncaptioned image] 10s 2K 92K 27.6
FunQA (Ours) Surprising Videos [Uncaptioned image] 19s 4K 312K 34.2
@@
Figure A1: Example of Magic Method Task. During the annotation process, we discovered that even as humans, it is difficult to fully understand the complete principles behind the implementation of magic tricks right from the beginning.
@@
Figure A2: FunQA Video Selection and Clip Principle. We have a zero-tolerance policy regarding the inclusion of offensive content in our dataset. During the video sourcing process (video selection and video clipping stages), we ensure that such content is completely eliminated.
@@
Figure A3: Examples of incorrect annotations and the suggested modifications. Here we clarify the meanings of different cell colors. Red: Errors made by the annotator during the annotation process. Yellow: Types of annotation errors. Green: Reasons for the errors and suggested corrections.
@@
Figure A4: GPT prompt for Translation and augmentation. Under our carefully designed prompt, GPT-3.5 can automatically generate English translations and five synonymous sentences from our high-quality Chinese annotations, thereby expanding FunQA dataset.
Figure A5: Issues in translation caused by Chinese-English cultural differences. In this example of translating and expanding a humorous video title, GPT-3.5 failed to understand the meaning of the original Chinese title, and we filtered out such data and made manual changes.
@@
Specifically for the Magic Method Task (M4), we found that most annotators were not professional magicians, and even when watching instructional magic videos, it was challenging for them to provide complete and reasonable annotations for magic methods. Therefore, for M4, we only provided a small number of examples, which do not exist in the FunQA dataset. Fig. A1 shows an example of M4.
We conducted strict quality reviews in real-time for the annotations, including but not limited to checking if the annotations meet the requirements, if the selected videos contain sensitive content, and if the annotation texts contain sensitive content. The annotation principles (video selection and video clip, etc) can be seen in Fig. A2. Fig. A3 also illustrates some errors encountered during the annotation process and the suggested modifications provided. We assured that every annotation included in the final release of FunQA has been subjected to rigorous multi-person, multi-round review processes. Specifically, every piece of released FunQA version annotated content underwent scrutiny by three distinct annotators. After completing all the annotations, we conducted a consensus evaluation study to assess the objectivity of our annotations. We randomly selected 10% of the videos and asked annotators to provide a consistency score (high consensus, medium consensus, low consensus) between their first and second annotations, considering all the previous annotations shown to them.
Copyright & License We respect the copyright of each video. Our data sharing approach is informed by the models established by previous studie like Kinetics [24], HD-VILA-100M [62], and others. Instead of providing the original raw data, we only supply the YouTube video IDs necessary for downloading the respective content.
@@
0.A.2 Language Augmentation
For each QA pair, we first asked GPT-3.5 to translate the previous Chinese answers into English, and then let GPT-3.5 give five separate answers with the same meaning but different linguistic expressions. Afterward, we filtered out incorrect and incomplete generations. Fig. A4 shows the prompt we gave to GPT-3.5.
However, as mentioned in the main text in the limitation, the translation of GPT-3.5 will be ambiguous when the original text uses Chinese harmonics and some special Chinese meanings, and this problem mostly occurs in the Title Task, we screened such problems by hand and modified them, as shown in Fig. A5.
0.A.3 Data Statistics
For a complete comparison between FunQA and other existing benchmarks, please see Table A1. We collected raw videos from multiple video platforms, including short videos, long videos, and video clips, and the detailed data can be seen in Table A2. The statics of FunQA dataset before GPT-3.5 extension can be seen in Table A3. The word cloud of the all annotation word is shown in Fig. 5(d) (a-d).
Table A2: Statistics of the FunQA raw data.
FunQA Subset Type Source # Videos Avg.len (s) Total.len (h)
Humor Daily Life (Human) Youtube 351 182 15.8
1296 14 5.18
Nature (Animal) 230 133 8.52
Creative Performance Youtube 26 6060 43.77
@@ 0.A.3 Data Statistics
Table A3: Statistics of the FunQA before GPT-3.5 extension.
Datasets Avg length (s) # Clips (K) # QA pairs (K) # QA per clip
HumorQA 7 1.8 7.2 4.0
CreativeQA 48 0.9 4.5 5.0
MagicQA 10 1.6 4.8 3.0
FunQA 19 4.3 16.5 3.8
(a) HumorQA.
(b) CreativeQA.
(c) MagicQA.
(d) FunQA.
Figure A6: Word cloud of FunQA free-text answer.
Appendix 0.B FunQA Extension Datasets
@@
Figure A7: GPT prompt for FunQA Multi-choice. We give GPT-3.5 prompt as system, and then enter the description and explanation of the video, and GPT-3.5 will output the data in the form of multiple-choice questions according to the prompt.
@@
Figure A8: GPT prompt for FunQA Dialog. We give GPT prompt as system, and then input the description and explanation of the video, and GPT-3.5 will output the data in the form of dialogue according to the prompt.
0.B.1 FunQA-MC Dataset
@@
FunQA-MC (Multi-choice) Dataset is prepared to provide training and testing for arbitrary models, in this dataset our QA pairs are in the form of multiple choice, the answer is a word, phrase, or short sentence, and the type of questions are all descriptions. An example of the prompt given to GPT-3.5 and the data generated can be seen in Fig. A7. For the FunQA-MC-R, which is a multi-choice version specifically containing counter-intuitive reasoning questions, we randomly chose 100 QA pairs from Task 2 ( Counter-intuitiveness Reasoning Task) to create the subset.
0.B.2 FunQA-DIA Dataset
@@
Most of the current LLMs are in the form of dialogues. To cater to their data input, we produced the FunQA-DIA (Dialog) dataset, in which we used GPT-3.5 to convert QA pairs into recursive dialogues with added context. In addition to this, we also let GPT-3.5 freely generate future dialogue development, which greatly expands the amount of information. An example of the prompt given to GPT-3.5 and the data generated can be seen in Fig. A8.
Appendix 0.C More Details of FunMentor
0.C.1 Real Fact Question
During the Real Fact Collection process, FunMentor relies on designing questions that VLMs can answer accurately and objectively to gain knowledge of the video’s basic information. The specific questions are shown in Fig. A9.
Figure A9: Example of Real fact question by FunMentor.
Figure A10: Prompt Design of FunMentor in multi-round dialogue. On the left are the prompts from FunMentor, which incorporate the information from collected Real Facts and conduct Answer Judgement to generate suggestions.
0.C.2 Prompts Design
Based on the QA pairs collected as mentioned, FunMentor engages in multi-round dialogues with VLMs through our predefined prompts. This process involves providing increasingly precise instructions, guiding the VLM towards correctly answering the question. Specific examples of this are shown in Fig. A10.
Table A4: Main Results on FunQA Benchmark (Complete Version). H2, C2, M2 represent the detailed video description task, and H3, C3, M3 represent reasoning around counter-intuitiveness. For the higher-level tasks, H4, C4 involve attributing a fitting and vivid title. The responses for all these tasks in free-text format. We use the following metrics: BLEU-4 / ROUGE-L / CIDEr (shown in the first row) and BLEURT / GPT-4 (shown in the second row) for evaluation. C5 represents scoring the video creativity, and the metric is the Accuracy between the predicted score and the official score. We tested the caption-based and instruction-based models. Here we clarify the abbreviation in the table. L.M.: GIT_LARGE_MSRVTT; L.V.: GIT_LARGE_VATEX; D.C. means finetuned on Dense Caption; FunQA means finetuned on FunQA.
HumorQA CreativeQA MagicQA
Task H2-Des. H3-Rea. H4-Title C2-Des. C3-Rea. C4-Title C5-Score M2-Des. M3-Rea.
- Caption-based Model
mPLUG [35] 1.5 / 16.4 / 1.0 19.9 / 3.9 1.1 / 12.5 / 0.4 25.7 / 6.0 0.6 / 7.5 / 0.1 22.1 / 11.2 0.4 / 13.4 / 0.0 14.9 / 3.0 0.7 / 12.6 / 0.1 24.2 / 6.9 0.3 / 3.2 / 0.0 20.8 / 18.8 0.0 1.2 / 15.8 / 0.5 19.7 / 4.0 0.9 / 8.9 / 0.4 21.2 / 8.1
GIT (L.M.) [57] 0.5 / 12.8 / 0.2 22.4 / 3.6 0.0 / 0.0 / 0.0 0.0 / 0.0 1.1 / 7.7 / 0.7 17.0 / 8.9 0.0 / 6.40 / 0.0 14.4 / 3.8 0.0 / 0.0 / 0.0 0.0 / 0.0 0.3 / 1.5 / 0.2 7.1 / 11.3 0.0 0.2 / 11.2 / 0.1 19.4 / 8.2 0.0 / 0.0 / 0.0 0.0 / 0.0
GIT (L.V.) [57] 1.2 / 16.9 / 0.6 33.3 / 4.0 0.0 / 0.0 / 0.0 0.0 / 0.0 1.0 / 8.8 / 0.7 25.9 / 10.0 0.1 / 8.30 / 0.0 20.5 / 4.2 0.0 / 0.0 / 0.0 0.0 / 0.0 0.5 / 2.8 / 0.4 10.5 / 12.0 0.0 0.6 / 13.7 / 0.1 29.8 / 8.6 0.0 / 0.0 / 0.0 0.0 / 0.0
- Instruction-based Model
VideoChat [36] 0.5 / 13.7 / 0.0 44.0 / 17.9 0.5 / 13.5 / 0.0 45.4 / 31.9 0.8 / 5.1 / 0.5 20.2 / 31.7 0.3 / 7.50 / 0.0 21.7 / 5.9 0.3 / 7.70 / 0.0 22.8 / 17.7 0.2 / 1.2 / 0.2 7.3 / 31.1 67.5 0.6 / 15.5 / 0.0 47.4 / 8.2 0.3 / 9.2 / 0.0 43.1 / 44.6
Video-ChatGPT [43] 0.5 / 14.0 / 0.1 39.9 / 24.3 0.7 / 12.4 / 0.1 40.1 / 24.9 0.4 / 3.2 / 0.2 36.5 / 41.2 1.1 / 19.8 / 0.2 45.8 / 6.6 0.8 / 17.3 / 0.1 45.2 / 9.1 0.2 / 1.9 / 0.2 30.9 / 48.8 85.4 0.7 / 20.8 / 0.0 50.8 / 11.2 0.5 / 11.3 / 0.0 43.3 / 40.4
mPLUG-Owl [65] 0.5 / 13.3 / 0.2 44.5 / 10,7 0.5 / 13.2 / 0.2 47.3 / 35.0 0.6 / 5.3 / 0.0 29.8 / 48.8 0.8 / 16.5 / 0.1 43.0 / 5.0 0.7 / 13.3 / 0.0 44.7 / 10.6 0.1 / 1.1 / 0.1 23.9 / 36.3 66.7 0.4 / 16.7 / 0.0 46.4 / 8.6 0.5 / 8.4 / 0.0 43.9 / 30.9
Video-LLaMA [70] 0.3 / 12.1 / 0.1 48.4 / 7.7 0.5 / 12.6 / 0.0 42.9 / 29.0 0.5 / 0.4 / 0.0 46.5 / 34.1 0.7 / 8.80 / 0.1 45.5 / 7.2 0.4 / 9.70 / 0.0 41.1 / 17.2 0.3 / 0.8 / 0.1 42.3 / 31.2 64.2 0.4 / 14.3 / 0.0 50.1 / 10.2 0.5 / 7.9 / 0.0 39.0 / 28.0
Otter (D.C.) [34] 1.1 / 14.3 / 0.4 30.2 / 7.7 1.2 / 14.2 / 0.4 32.3 / 28.3 0.5 / 5.4 / 0.1 21.7 / 20.0 0.5 / 13.8 / 0.1 28.7 / 1.7 1.0 / 16.8 / 0.2 32.9 / 7.9 0.3 / 2.3 / 0.1 17.7 / 36.3 45.0 1.0 / 15.0 / 0.3 32.5 / 2.1 1.1 / 12.8 / 0.2 27.3 / 36.8
Otter (FunQA) [34] 1.5 / 18.1 / 0.9 38.4 / 8.9 1.3 / 15.4 / 0.5 42.6 / 31.7 0.8 / 5.9 / 0.5 47.5 / 32.1 1.5 / 19.6 / 0.5 40.0 / 7.3 2.2 / 21.2 / 0.5 41.1 / 8.8 0.3 / 4.3 / 0.3 44.5 / 38.8 69.4 2.6 / 23.8 / 1.6 44.7 / 10.3 3.4 / 20.3 / 2.6 44.5 / 47.5
Video-ChatGPT [43] + FunMentor (Ours) 0.6 / 14.2 / 0.6 65.2 / 33.2 1.1 / 14.4 / 0.3 57.5 / 36.5 1.0 / 5.7 / 0.6 50.2 / 65.1 1.4 / 20.2 / 0.3 66.3 / 14.2 1.4 / 18.8 / 0.2 58.7 / 23.4 0.3 / 2.2 / 0.3 45.3 / 52.2 85.4 1.1 / 23.2 / 0.4 55.1 / 13.3 2.1 / 14.4 / 0.9 46.3 / 54.8
Otter (FunQA) [34] + FunMentor (Ours) 1.9 / 20.2 / 1.1 33.4 / 13.4 1.7 / 19.3 / 0.9 37.8 / 45.8 0.9 / 6.8 / 0.9 58.3 / 34.2 2.4 / 23.1 / 1.1 60.4 / 11.0 3.3 / 26.5 / 0.7 44.4 / 9.3 0.4 / 4.7 / 0.3 53.9 / 43.5 69.4 3.1 / 25.5 / 2.5 43.5 / 12.81 5.6 / 21.2 / 3.4 38.91 / 56.4
Appendix 0.D More Details of Experiment
The complete result of FunQA Benchmark (with traditional metircs’ scores) are shown in Table A4.
0.D.1 Significance of New GPT-4 Based Metric